Calculate the total area of the region described. Do not count area beneath the x-axis as negative.
Bounded by the curve y = 7(sqrt.x), the x-axis, and the lines x = 0 and x = 4

Answer by Alan3354(69443)

Same problem, different numbers.
-------------------------------------
Calculate the total area of the region described.
Bounded by the curve y=sqrt(x), the x-axis, and the lines x = 0 and x = 1
----
y = x^(1/2)
INT = F(x) = x^(3/2)/(3/2) = 2x^(3/2)/3 --- ignore the constant of integration.
F(1) = 2/3
F(0) = 0
Area = 2/3 sq units